midpoint of the line segment with the given endpoints.
(2,6) and (7,3)
thank you

Answer by stanbon(75887)

midpoint of the line segment with the given endpoints.
(2,6) and (7,3)
---------------------
Let (x,y) be the midpoint.
---
x = (2+7)/2 = 4.5
y = (6+3)/2 = 4.5
========================
midpoint = (4.5 , 4.5)
=============================
Cheers,
Stan H.